A volume of 1,5 m3 is occupied by two fluids. Fluid 1 has a specific gravity of SG=2, while Fluid 2 has a density of 800 kg.m-3? The mass of Fluid 1 is three times the mass of Fluid 2. calculate each fluids mass


1
Expert's answer
2021-10-10T16:02:30-0400

Express the volume in terms of mass and density:


"V=\frac{3m}{\rho_1}+\frac{m}{\rho_2},\\\space\\
m=V/\bigg(\frac{3}{\rho_1}+\frac{1}{\rho_2}\bigg)=545\text{ kg}."

Therefore, the mass of the first fluid is three times greater, or 1636 kg.